• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Holmstrom

    Log entry time 09:37:44 on May30,2007

    Entry number 203662

    keyword=Move to Q2=34

    We have finished Q2=24 and have started to move Q2=34.

    We have changed the left spectrometer to 1.895 GeV and 33.2 deg

    We have changed the right spectrometer to 0.5389 GeV and 65.046 deg

    We have entered 1895 and 1078 into the trigger download

    We have changed the Right arm S1s to 73% voltage.

    We are waiting ont he right dipole

    FIGURE 1

    FIGURE 2